OpenStudy (anonymous):

How many ways are there to choose a committee of 7 people from a group of 10 people?

OpenStudy (anonymous):

you are being asked to compute "10 choose 7" sometimes written as \[\dbinom{10}{7}\] or \[_{10}C_7\] do you know how to compute it?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

answers are 720,120,840,and 5040

OpenStudy (anonymous):

how many diffrent ways can you choose 7 people out of 10? is that what its asking.

OpenStudy (anonymous):

yes. answers are 720,120,840,and 5040

OpenStudy (anonymous):

Permutations and Combinations

OpenStudy (anonymous):

In how many ways can a committee of 3 people be chosen from 20 people? Since it's a "committee" (no mention of chairman or any other named position), the order of choosing the members is not important. Hence, you need to use the *combinations* rather than *permutations* formula. 20 choose 3 = 20! / (17! 3!) = 20 * 19 * 18 / 3! = 1140 52 choose 3 = 52! / (49! 3!) = 52 * 51 * 50 / 3! = 2210 So try to apply this to your question.

OpenStudy (anonymous):

CANT YOU JUST GIVE ME THE ANSWER :(

OpenStudy (anonymous):

i want to say its 840 but im not 100% sure so i think you should double check!